Translate the sentence to an equation. Then solve the equation. Three times the difference of some number and 6 amounts to the quotient of 99 and 11. Find the number. Write the equation. Use x to represent a number. Choose the correct answer below. ( A. 3 99 x-6 11 OB. ? ?. 3(x-6)-99-11 O D 3(x + 6) =- The number is
Answer #1

1. Three times the difference of number and 6.

let the number be x

Split this

1a) difference of number and 6 = x-6

1b)three times of it. = 3(x-6)

2. quotient of 99 and 11= 99/11

So, the translation would be,

3(x-6) = 99/11

Option C

further if we solve for x in the above equation,

we get, x-6 = 3 and value of x = 9